40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 230     Entry time: Tue Jun 18 19:05:16 2019
Author: Jon 
Type: Lab Infrastructure 
Category: General 
Subject: Large optical table cleared and set up for point absorbers 

Aidan and I continued the lab clean-up today. There's still more to do, but we did fully clear the large optical table which formerly housed the 50 W CO2 laser. I moved the optical enclosure over from the small table to serve as the area for the point absorber experiment. Inside it I mounted the Hartmann sensor and a 532 nm Thorlabs LED source. The LED still needs collimating/focusing optics to be installed.

Attachment 1: IMG_3421.jpg  2.120 MB  Uploaded Tue Jun 18 20:08:03 2019  | Hide | Hide all
IMG_3421.jpg
ELOG V3.1.3-